OpenStudy (anonymous):

Find an equation of the sphere with center (4, 4, 1) and radius 4.

OpenStudy (anonymous):

let center = (x1, y1, z1) = (4,4,1) equation of a sphere: (x-x1)^2 + (y-y1)^2 + (z-z1)^2 = r^2 (x-4)^2 + (y-4)^2 +(z-1)^2 = 16

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

thats what i got too...just wasnt sure if it was correct

OpenStudy (anonymous):

no problem ; )